# Towards More Unified In-context Visual Understanding

### 05 Dec 2023 v1

@@ 1 Introduction
More recently, some works start to apply the similar ICL idea into the vision-only tasks via formulating the learning goal as image inpainting [4, 46, 48]. With the well-collected multi-task vision datasets and unified grid image prompt design, these works utilize pre-trained masked image modeling models to give a perspective of what can be general-purpose task prompts in vision. For instance, SegGPT [48] studies the fundamental visual understanding problem, segmentation task, as an in-context coloring problem to achieve the in-context segmentation capability. Yet, the pre-trained vision-centric inpainting framework confines the output modality to be image only. Therefore, a straightforward question is “How to perform in-context learning with multimodal output enabled for visual understanding in a unified framework?”
@@ 2 Related Works
Within the field of computer vision, the study [4] initially advances an in-context training paradigm utilizing image inpainting on illustrations and infographics derived from vision-related literature, which shows competencies in fundamental CV tasks such as foreground segmentation, singular object detection, and colorization. Additionally, the study by Painter [46] employs masked image modeling on continuous pixels to conduct in-context training with self-organized supervised datasets in seven tasks, and yields highly competitive outcomes on them. Subsequently, SegGPT [48] is a dedicated method trying to solve diverse and unlimited segmentation tasks with a similar framework. Recent studies have concentrated on how to enhance the ICL capability in vision, such as prompt selection [41] and the execution of nearest neighbor retrieval utilizing a memory bank [3].
@@ 3.1 Vision-Language Prompt Design
We begin by implementing unified vision-language prompts to depict different types of vision-language tasks. We treat $k$ in-context samples with input and output like ”$(i_{1},o_{1}),\cdots,(i_{k+1},o_{k+1})$” as interleaved data, and embed them in the discrete token space. This innovative design provides the flexibility required for customizing vision or vision-language tasks according to specific needs and preferences.
Vision-Only Tasks. Following previous works, we conduct all vision-only tasks as an inpainting task. However, the inpainting is performed in token space. For every image pair that is composed of an original image and its corresponding task output, we first quantize them into discrete tokens utilizing a pre-trained image quantizer. A special tag “[BOI]” is inserted in front of each image’s token representation. Then we concatenate each pair’s visual tokens obeying the order of precedence. This structure creates a cohesive relationship between the two in-context pairs, framing them both as visual token components.
Vision-Language Tasks. For vision-language tasks, here we take the dense captioning task as an example. The prompts are clear and closely resemble those of natural language processing (NLP) tasks. Similar to existing methods [1], multiple captioning samples can be treated as interleaved image and text data. For each image, we quantize them the same way as in vision-only tasks, with the special “[BOI]” tag. For the text part, we describe the region caption with their corresponding instance category and bounding box (bbox) like “Category: <$\textless c\textgreater$ >. Bboxes: [$x_{1},y_{1},x_{2},y_{2}$]. Caption: <$\textless text\textgreater$ >.” While $c$ denotes the class index in the category set, and $P=\{x_{i},y_{i}\}^{N}_{i=1}$ represents points that locate the object. <$\textless text\textgreater$ > represents the placeholder of caption tokens. We also add a special tag “[BOT]” at the beginning of each caption. After being tokenized by looking up the vocabulary, we use a similar concatenation strategy to get the in-context token representations, which streamlines the understanding and implementation of vision-language tasks.
At the conclusion of each segment of in-context tokens, we incorporate an “[EOC]” tag to signify the completion of in-context samples.
@@
For the text part, the subword Byte-Pair Encoding (BPE) tokenizer in GPT-2 [34] is utilized. Given a text $x_{text}$, by subwords frequency analysis and iteratively merging most common adjacent tokens, we update the vocabulary as well as the text with merged tokens until the termination condition. Then in the context of encoding information, BPE tokenizer quantizes $x_{text}$ into tokens $z_{q\_t}$ by looking up the vocabulary.
@@ 4 Experiments
For the image tokenizer, we adopt VQ-GAN tokenizer [13] with a vocabulary size of 1024 and 16x downsampling ratio, which is pre-trained on the Imagenet dataset. The input image resolution is set to $256\times 256$, leading to 256 tokens after quantization. For the text tokenizer, we employ GPT-2 BPE tokenizer [34] with a vocabulary size of 50257. We implement our model with GPT-small model architecture while replacing the FFN in part of the decoder layers with attribute routing MoEs introduced in [58]. Please refer to the supplementary for detailed architecture hyperparameters. For MoE layers, top-2 gating is utilized for gate decisions, and the number of experts in each MoE layer is set to 8 by default.
@@ 4.3 Ablation Studies
Class-aware In-context Task Definitions. In our exploration of two proposed in-context learning tasks, we rigorously examine the task definitions. As demonstrated in Table 1, we investigate the object size and scale within each in-context sample for the CA-ICL segmentation task. The term ”diverse sizes” refers to the decision on whether to exclude small instances, while the ”large scale” setting denotes the choice between random shift cropping based on the mask region or otherwise using short edge cropping in image processing. Our findings indicate that including small objects with a large object scale yields optimal results. We surmise that objects spanning multiple scales offer more detailed insights and salient in-context samples lead to a richer diversity of information, which is beneficial for segmentation.
@@
In our research on CA-ICL captioning, We explore the correlation between in-context input images and their corresponding descriptions. We drew inspiration from dense captioning and visual grounding, examining if incorporating object location information is beneficial for the model to capture semantic cues conveyed by in-context samples. We experimented with two distinct bounding box definitions. The first method requires the model to output bbox of the target object as a binary mask where zero represents the target, and one signifies the background. The second method integrates bbox information directly into the caption, as elaborated in Section 3.1.
@@
As evidenced in Table 2, introducing an image-type output leads to a notable decline in performance compared to the baseline. This might be attributed to the fact that the output bbox image contains more tokens than the caption (64 vs. typically fewer than 20), therefore the model mainly focuses on modeling the mask part. To tackle this issue, we explored the method of encoding bbox information in a textual format, as outlined in Section 3.1. While the results were considerably better than the “bbox_image” approach, even outperformed the baseline in CIDEr metric. Figure 5 demonstrates that using prompts of the “bbox_text” type leads to more precise predicted captions that correspond with the intended region. This alignment significantly aids in the accurate and convenient verification of the model’s performance during testing phases. This evidence supports the model’s capability to effectively generate class-aware captions when supplied with appropriate examples.
@@
Model Variants Definition. We conducted experiments using various model configurations at a higher resolution of 256, maintaining constant other hyperparameters as discussed in the task definition paragraph, to identify the optimal choice. The reference for these experiments is the single task performance, with the baseline established as task co-training using the standard GPT-2 small architecture, referred to as “all tasks”. We replace the FFN in part of transformer blocks with the MoE layer proposed in [23] and the AG_MoE introduced in [58] for analysis. The results presented in Table 3 reveal that the baseline setting results in significant unbalanced performance with a sharp segmentation performance decrease, while models with MoE configurations surpass the baseline in segmentation performance by 18.74 scores, yet there remains a notable shortfall of 10.8 scores in captioning performance. The adoption of the AG_MoE structure further narrows this performance gap. We surmise that this discrepancy may be due to the disproportionate length of image tokens relative to text tokens. Models with shared parameters might struggle to effectively manage the significant difference in token representations between the two tasks, highlighting the advantages of MoEs. In the following section, we will address the challenges associated with multi-task co-training.
@@
To further enhance the performance of the model with AG_MoE, we adopt a multi-task learning paradigm to alleviate the task interference problems and, meanwhile, stabilize the training of MoEs. Drawing inspiration from Uni-Perceiver v2 [25], we utilize their unmixed batch sampling strategy and correlative optimizer. Here, the sampling weight $s_{k}$ of each dataset is configured to be proportional to the square root of the dataset’s size. For the scaling factor $w_{k}$, we uniformly assign a value of 1 to all tasks. As evidenced in Table 3, the integration of the AG_MoE architecture with our multi-task learning strategy results in performance that exceeds the baseline for both tasks. This is particularly notable in the CA-ICL segmentation task, where an impressive gain of 28.17 points is observed. This indicates that the multi-task strategy effectively prevents potential task conflicts within a batch through the task non-overlapping sampling mechanism as well as gradient normalization.
@@ 4.4 Comparison with State-of-the-art Methods
In the domain of CA-ICL captioning, the generalist baseline for evaluation is Openflamingo [2], a large vision-language model that excels in demonstrating strong in-context captioning ability. The CA-ICL captioning task most analogous to it is that of dense captioning, as both tasks necessitate the prediction of not only the caption but also the corresponding bbox. Therefore, we compare with the sota dense captioning model GRiT [50]. We utilize the images in our test set to evaluate GRiT. Then allocate the generated predictions to our ground-truth regions annotations, utilizing the IoU metric of their respective bboxes as the basis for the assignment. As shown in Table 4, our method achieves state-of-the-art performance in traditional image captioning metrics. In comparison to Openflamingo, which has a parameter tenfold greater, we also achieve a 0.7-point increase in BLEU4 and a significant 25.6-point improvement in CIDEr. However, the result still has a gap in the mAP score compared with GRiT. We believe this is due to the fact that they utilize a foreground object extractor to locate the target.
@@ 5 Conclusion
In this work, we present a unified framework for in-context visual understanding. By leveraging multimodal quantization and unified embedding, our model is capable of jointly learning multimodal data in the general token embedding space. The interleaved prompts enable the in-context ability and facilitates advanced semantic understanding. By synergistically integrating autoregressive transformer with the MoEs framework, we achieve stable multi-task co-training while simultaneously benefiting from the balanced contributions of each task. Based on the class-aware in-context benchmark, extensive experiments demonstrate the effectiveness and generality of our model. Overall, our research showcases the potential of in-context learning across various modalities as well as tasks.
Appendix A Model Architecture and Configuration
@@
Appendix B Additional Analysis
Loss and padding analysis for CA-ICL captioning. In this section, we delve into the effects of loss functions and padding strategies on captioning performance. We employ the baseline loss $\mathcal{L}_{out}$, delineated in Section 3.3 of our main submission, aligning the length of caption tokens with image tokens via padding. We examine different CE loss weights for the input image tokens in each context (denoted as $\mathcal{L}_{in}$). This setting is based on the intuition that image captioning task may benefit from a increased focus on visual content because of the unbalanced sequence length between the image and text tokens. As indicated in Table 6, the experiment results reveal that using only $\mathcal{L}_{out}$ surpasses other configurations with $\mathcal{L}_{in}$ in a one-shot setting. In contrast, a composite loss of $0.5\mathcal{L}_{in}+\mathcal{L}_{out}$ achieves superior results in two- and three-shot scenarios. Consequently, we adopt the $0.5\mathcal{L}{in}+\mathcal{L}{out}$ loss for individual captioning tasks, while utilizing a consistent $\mathcal{L}_{out}$ during co-training sessions.
In-context effectiveness analysis. We study the impact of increasing the number of in-context pair examples in the task prompt on the outcomes. We trained the model using three in-context samples and inference with 1 to 3 samples. As presented in Table 7, for the CA-ICL segmentation task, the inclusion of more examples consistently improves the segmentation performance. However, for the CA-ICL captioning task, the performance does not exhibit a steady increase. The possible reason is that using more in-text samples for the segmentation task can provide more segmentation clues coming from different views, appearances of different image samples for the same category of target object. But for the caption task, one caption is already enough to denote the target object while multiple description styles from different samples will introduce more style ambiguity.
@@ Appendix B Additional Analysis
Appendix C Additional Qualitative Results
@@
Appendix D Limitation
Because of the long-tailed class and object scale distribution of the training dataset, the model does not perform well with multiple small objects or uncommon classes like traffic light. Some typical failure cases are presented in Figure 10 and Figure 11. We think a more balanced data distribution may be beneficial for the situation. For example, utilizing Copy-Paste strategies [12, 56] to expand the per-category instance pool. For improving captioning, a potential solution involves resampling the data or other data balancing strategies. We leave this issue to future research.

### 16 Mar 2024 v2

@@ 1 Introduction
More recently, some works start to apply the similar ICL idea into the vision-only tasks via formulating the learning goal as image inpainting [4, 47, 48]. With the well-collected multi-task vision datasets and unified grid image prompt design, these works utilize pre-trained masked image modeling models to give a perspective of what can be general-purpose task prompts in vision. For instance, SegGPT [48] studies the fundamental visual understanding problem, segmentation task, as an in-context coloring problem to achieve the in-context segmentation capability. Yet, the pre-trained vision-centric inpainting framework confines the output modality to be image only. Therefore, a straightforward question is “How to perform in-context learning with multimodal output enabled for visual understanding in a unified framework?"
@@ 2 Related Works
Within the field of computer vision, the study [4] initially advances an in-context training paradigm utilizing image inpainting on illustrations and infographics derived from vision-related literature, which shows competencies in fundamental CV tasks. Additionally, the study by Painter [47] employs masked image modeling on continuous pixels to conduct in-context training with self-organized supervised datasets in seven tasks, and yields highly competitive outcomes on them. Subsequently, SegGPT [48] is a dedicated method trying to solve diverse and unlimited segmentation tasks with a similar framework. Recent studies have concentrated on how to enhance the ICL capability in vision, such as prompt selection [41] and the execution of nearest neighbor retrieval utilizing a memory bank [3].
@@ 3.1 Vision-Language Prompt Design
We begin by implementing unified vision-language prompts to depict different types of vision-language tasks. We treat $k$ in-context samples with input and output like "$(i_{1},o_{1}),\cdots,(i_{k+1},o_{k+1})$" as interleaved data, and embed them in the discrete token space. This innovative design provides the flexibility required for customizing vision or vision-language tasks according to specific needs and preferences.
Vision-Only Tasks. Following previous works, we conduct all vision-only tasks as an inpainting task. However, the inpainting is performed in token space. For every image pair that is composed of an original image and its corresponding task output, we first quantize them into discrete tokens utilizing a pre-trained image quantizer. A special tag “[BOI]" is inserted in front of each image’s token representation. Then we concatenate each pair’s visual tokens obeying the order of precedence. This structure creates a cohesive relationship between the two in-context pairs, framing them both as visual token components.
Vision-Language Tasks. For vision-language tasks, here we take the dense captioning task as an example. The prompts are clear and closely resemble those of natural language processing (NLP) tasks. Similar to existing methods [1], multiple captioning samples can be treated as interleaved image and text data. For each image, we quantize them the same way as in vision-only tasks, with the special “[BOI]" tag. For the text part, we describe the region caption with corresponding instance category and bounding box (bbox) like “Category: <$\textless c\textgreater$ >. Bboxes: [$x_{1},y_{1},x_{2},y_{2}$]. Caption: <$\textless text\textgreater$ >." While $P=\{x_{i},y_{i}\}^{N}_{i=1}$ represents points that locate the object. <$\textless text\textgreater$ > represents the placeholder of caption tokens. We also add a special tag “[BOT]" at the beginning of each caption. After being tokenized by looking up the vocabulary, we use a similar concatenation strategy to get the in-context token representations.
At the conclusion of each segment of in-context tokens, we incorporate an “[EOC]" tag to signify the completion of in-context samples.
@@ 4 Experiments
For the image tokenizer, we adopt VQ-GAN tokenizer [13] with a vocabulary size of 1024 and 16x downsampling ratio, which is pre-trained on the Imagenet dataset. The input image resolution is set to $256\times 256$, leading to 256 tokens after quantization. For the text tokenizer, we employ GPT-2 BPE tokenizer [34] with a vocabulary size of 50257. We implement our model with GPT-small model architecture while replacing the FFN in part of the decoder layers with attribute routing MoEs introduced in [58]. Please refer to the supplementary for detailed architecture hyperparameters.
@@ 4.3 Ablation Studies
Class-aware In-context Task Definitions. In our exploration of two proposed in-context learning tasks, we rigorously examine the task definitions. As demonstrated in Table 1, we investigate the object size and scale within each in-context sample for the CA-ICL segmentation task. Our findings indicate that including small objects with a large object scale yields optimal results. We surmise that objects spanning multiple scales offer more detailed insights and salient in-context samples lead to a richer diversity of information, which is beneficial for segmentation.
@@
In our research on CA-ICL captioning, We explore the correlation between in-context input images and their corresponding descriptions. We drew inspiration from dense captioning and visual grounding, examining if incorporating object location information is beneficial for the model to capture semantic cues conveyed by in-context samples.
@@
As evidenced in Table 2, introducing an image-type output leads to a notable decline in performance compared to the baseline. To tackle this issue, we explored the method of encoding bbox information in a textual format, as outlined in Section 3.1. While the results were considerably better than the “bbox_image" approach, even outperformed the baseline in CIDEr metric. Figure 5 demonstrates that using prompts of the “bbox_text" type leads to more precise predicted captions that correspond with the intended region. This alignment significantly aids in the accurate and convenient verification of the model’s performance during testing phases. This evidence supports the model’s capability to effectively generate class-aware captions when supplied with appropriate examples.
@@
Model Variants Definition. We conducted experiments using various model configurations at a higher resolution of 256 to identify the optimal choice. The reference for these experiments is the single task performance, with the baseline established as task co-training using the standard GPT-2 small architecture, referred to as “all tasks". We replace the FFN in part of transformer blocks with the MoE layer proposed in [23] and the AG_MoE introduced in [58] for analysis. The results presented in Table 3 reveal that the baseline setting results in significant unbalanced performance with a sharp segmentation performance decrease, while models with MoE configurations surpass the baseline in segmentation performance by 18.74 scores, yet there remains a notable shortfall of 10.8 scores in captioning performance. The adoption of the AG_MoE structure further narrows this performance gap. Considering the image tokens dominate compared with text tokens and the divergent gradient directions of differing task complexities (as shown in Figure 6), the caption performance drops. Models with shared parameters might struggle to effectively manage the significant difference in token representations between the two tasks, highlighting the advantages of MoEs. In the following section, we will address the challenges associated with multi-task co-training.
@@
Figure 6: Gradient comparison of CA-ICL tasks. We utilize the normalized average gradient of each attention layer for comparison, while the symbol and the value represent the direction and magnitude of the gradient respectively.
@@
To further enhance the performance of the model with AG_MoE, we adopt a multi-task learning paradigm to alleviate the task interference problems and, meanwhile, stabilize the training of MoEs. Drawing inspiration from Uni-Perceiver v2 [25], we utilize their unmixed batch sampling strategy and correlative optimizer. Here, the sampling weight $s_{k}$ of each dataset is configured to be proportional to the square root of the dataset’s size. For the scaling factor $w_{k}$, we uniformly assign a value of 1 to all tasks. As evidenced in Table 3, the integration of the AG_MoE architecture with our multi-task learning strategy results in performance that exceeds the baseline for both tasks. This is particularly notable in the CA-ICL segmentation task, where an impressive gain of 28.17 points is observed. This indicates that the multi-task strategy effectively prevents potential task conflicts within a batch.
@@ 4.4 Comparison with State-of-the-art Methods
In the domain of CA-ICL captioning, the generalist baseline for evaluation is Openflamingo [2], a large vision-language model that excels in demonstrating strong in-context captioning ability. The CA-ICL captioning task most analogous to it is that of dense captioning, as both tasks necessitate the prediction of not only the caption but also the corresponding bbox. Therefore, we compare with the sota dense captioning model GRiT [50]. We utilize the images in our test set to evaluate GRiT. Then allocate the generated predictions to our ground-truth regions annotations, utilizing the IoU metric of their respective bboxes as the basis for the assignment. As shown in Table 4, our method achieves state-of-the-art performance in traditional image captioning metrics. In comparison to Openflamingo, which has a parameter tenfold greater, we also achieve a 0.7-point increase in BLEU4 and a significant 25.6-point improvement in CIDEr. However, the result still has a gap in the mAP score compared with GRiT. We believe this is because they utilize a foreground object extractor.
@@ 5 Conclusion
In this work, we present a unified framework for in-context visual understanding. By leveraging multimodal quantization and unified embedding, our model is capable of jointly learning multimodal data in the general token embedding space. By synergistically integrating autoregressive transformer with the MoEs framework, we achieve stable multi-task co-training while simultaneously benefiting from the balanced contributions of each task. Overall, our research showcases the potential of in-context learning across various modalities as well as tasks.
[…]
Supplementary Material
Hyperparameter GPT-2 Ours
Architecture transformer decoder transformer decoder
Vocabulary size 50257 51290 / 52295
Max positions 1024 2060
Hidden size 768 768
Hidden layers 12 12
Attention heads 12 12
Number of MoEs - 6
Number of experts - 8
Layer norm epsilon 1e-05 1e-12
Attention probs dropout prob 0.1 0.1
Hidden dropout prob 0.1 0.1
Table 5: Hyperparameters for our GPT-2 baseline and proposed model. Note for the vocabulary size, we experiment with two settings for whether to add the special category and bbox tokens.
B@4 $\uparrow$ CIDEr $\uparrow$
Examples 1 2 3 1 2 3
$\mathcal{L}_{out}$ 5.4 1.8 1.6 95.9 52.5 51.2
w/ $\mathcal{L}_{in}$ 4.5 1.9 1.9 85.0 54.7 54.5
w/ $0.5\mathcal{L}_{in}$ 5.3 2.0 2.0 86.6 54.6 55.6
Table 6: Loss analysis on class-aware in-context captioning task.
Examples CA-ICL segmentation CA-ICL captioning
MIoU normal-$\uparrow$ MAE normal-$\downarrow$ B@4 normal-$\uparrow$ CIDEr normal-$\uparrow$ mAP normal-$\uparrow$
0 45.70 0.094 2.8 65.7 0.2
w/o category information for CA-ICL captioning
1 56.17 0.167 6.6 95.6 1.5
2 59.21 0.132 1.7 45.5 1.7
3 60.85 0.128 0.8 32.6 1.6
w category information for CA-ICL captioning
1 58.04 0.110 5.3 86.9 10.9
2 61.65 0.101 2.3 60.9 0.8
3 62.33 0.098 2.3 62.2 1.5
Table 7: Analysis on in-context samples and category information. We report the metrics utilized in our main experiments for the two CA-ICL tasks.
6 Model Architecture and Configuration
@@
7 Additional Quantative Analysis
Loss and padding analysis for CA-ICL captioning. In this section, we delve into the effects of loss functions and padding strategies on captioning performance. We employ the baseline loss $\mathcal{L}_{out}$, delineated in Section 3.3 of our main submission, aligning the length of caption tokens with image tokens via padding. We examine different CE loss weights for the input image tokens in each context (denoted as $\mathcal{L}_{in}$). This setting is based on the intuition that image captioning task may benefit from an increased focus on visual content because of the unbalanced sequence length between the image and text tokens. As indicated in Table 6, the experiment results reveal that using only $\mathcal{L}_{out}$ surpasses other configurations with $\mathcal{L}_{in}$ in a one-shot setting. In contrast, a composite loss of $0.5\mathcal{L}_{in}+\mathcal{L}_{out}$ achieves superior results in two- and three-shot scenarios. Consequently, we adopt the $0.5\mathcal{L}{in}+\mathcal{L}{out}$ loss for individual captioning tasks, while utilizing a consistent $\mathcal{L}_{out}$ during co-training sessions.
In-context effectiveness analysis. We study the impact of increasing the number of in-context pair examples and whether to add category information in CA-ICL captioning in the task prompt on the outcomes. We trained the model using three in-context samples and inference with 1 to 3 samples. Additionally, we trained one model with explicit class information input instead of in-context samples for comparison. As presented in Table 7, with only class information, the model performs pool on both tasks for the CA-ICL segmentation task, which indicates the effectiveness of in-context samples as they are given more information than the simple class label. The inclusion of more examples consistently improves the segmentation performance. However, for the CA-ICL captioning task, the performance does not exhibit a steady increase, even more serious if category information is not provided. The possible reason is that using more in-text samples for the segmentation task can provide more segmentation clues coming from different views, and appearances of different image samples for the same category of target object. But for the caption task, one caption is already enough to denote the target object while multiple description styles from different samples will introduce more style ambiguity. From the perspective of performance, we report the best results of the model with class information for the captioning task. The problem of captioning is left to further study.
Time cost of different models. We calculate the fps metric for the 1 in-context example setting to analyze the time cost. The inference speed of our model using 0, 1, 2, 3 in-context examples is 2.8, 2.4, 1.9, 1.4 fps, respectively. While for SegGPT and OpenFlamingo, the fps is 7.7 and 0.3 img/s. Our model is capable of using in interactive applications.
@@ 7 Additional Quantative Analysis
8 Additional Qualitative Results
@@
9 Limitation and future work
Because of the long-tailed class and object scale distribution of the training dataset, the model does not perform well with multiple small objects or uncommon classes like traffic light. Some typical failure cases are presented in Figure 11 and Figure 12. We think a more balanced data distribution may be beneficial for the situation. For example, utilizing Copy-Paste strategies [12, 56] to expand the per-category instance pool. For improving captioning, a potential solution involves resampling the data or other data balancing strategies. Another limitation is that the model only supports one class per forward. Currently, the proposed model can support multiple categories by multiple times inference. The color mapping strategy utilized in SegGPT might be helpful.
[…]